OpenStudy (anonymous):

A rectangular box has a square base with an edge length of x cm and a height of h cm. The volume of the box is given by V = x^2h cm^3. Find the rate at which the volume of the box is changing when the edge length of the base is 12 cm, the edge length of the base is decreasing at a rate of 2 cm/min, the height of the box is 6 cm, and the height is increasing at a rate of 1 cm/min. @freckles

OpenStudy (anonymous):

The volume of the box is decreasing at a rate of 144 cm^3/min. The volume of the box is increasing at a rate of 288 cm^3/min. The volume of the box is decreasing at a rate of 288 cm^3/min. The volume of the box is increasing at a rate of 144 cm^3/min.

OpenStudy (freckles):

did you differentiate the equation given w.r.t to time variable t ?

OpenStudy (freckles):

\[V(t)=[x(t)]^2 \cdot h(t)\] use product rule and chain rule

OpenStudy (anonymous):

\[V(t)=2(u)(1)(dh(t)/dx)+h(t) (2)\]

OpenStudy (anonymous):

okay I got the same thing

OpenStudy (freckles):

\[V(t)=x^2h \\ V'(t)=2x x' h+x^2h'\] is this what you mean?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[V'=? \\ x=12 \\ x'=-2 \\ h=6 \\ h'=1 \] just enter in the values

OpenStudy (freckles):

and evaluate V'

OpenStudy (freckles):

\[V'=2x x' h+x^2h' \\ V'=2(12)(-2)(6)+(12)^2(1)\] All I did was just replace the values in the equation with the values I gave me

OpenStudy (freckles):

just use order of operations

OpenStudy (anonymous):

V'=-144 @freckles

OpenStudy (anonymous):

so the answer is A

OpenStudy (freckles):

24(-12)+144 12(-24+12) 12(-12) -144 yes seems right the negative there means the volume is decreasing and it is decreasing at a rate of 144 cm^3/min

OpenStudy (anonymous):

ok great thanks so much!